C = 120 + 0.75 Yd,

I = 250 + 0.3 Y - 5i,

T = 50 + 0.15Y,

G = 900,

X = 550,

M = 90 + 0.3Y,

Md = 8Y - 2i,

Ms = 2000,



C = consumption,

Yd - disposable income,

I – investment spending,

i - interest rate,

T - tax income,

G – government spending,

X – exports,

M - imports,

Md - money demand and

Ms - Money supply.


Please derive the IS and LM Equation from the above information.





1
Expert's answer
2020-11-20T07:23:14-0500

For IS


"i=\frac {120+250+0.3Y+460}{5}-\frac{1-0.75+0.3}{5}Y+\frac{1}{5}G-\frac{0.75}{5}T_a"


"Y=\frac{120+250+460}{1-0.75+0.3}+\frac{1}{0.55}G-\frac{0.75}{0.55}T_a"


For LM

"i=\frac{8}{-2}Y-\frac{1}{-2}M_d"

"Y=\frac{1}{8}M_d+\frac{-2}{8}i"